2/3(x-6)+x=3x-2 wat does x equal please shpw work i dont get it

Answer by rfer(16322)

2/3(x-6)+x=3x-2
2/3x-4+x=3x-2
1 2/3x=3x+2
1 1/3x=2
4/3x=6/3
x=(6/3)/(4/3)
x=6/3*3/4
x=18/12
x=1 1/2